Measures of liquidity, Solvency, and Profitability

The comparative financial statements of Marshall Inc. are as follows. The market price of Marshall common stock was $ 59 on December 31, 20Y2

Determine the following measures for 20Y2, rounding to one decimal place, except for dollar amounts, which should be rounded to the nearest cent. Use the rounded answer of the requirement for subsequent requirements, if required. Assume 365 days a year

see screenshots, please. I need help solving 4-18. Measures of liquidity, Solvency, and Profitability The comparative financial statements of Marshall Inc. are as follows. TheMarshall Inc. Comparative Income Statement For the Years Ended December 31, 20Y2 and 20Y1 20Y2 20Y1 Sales $2,052,060 $2,227,2Marshall Inc. Comparative Balance Sheet December 31, 20Y2 and 2041 20Y2 20Y1 Assets Current assets Cash $421,970 $367,450 638Required: Determine the following measures for 20Y2, rounding to one decimal place, except for dollar amounts, which should bThanks
Answer #1

Answer:

4.Accounts Receivable Turnover = Sales / Avg. Accounts Receivable = 2227230 / ( 423400 + 401500 ) / 2

                                                                                                    = 2227230 / 412450 = 5.4

5.Number of Days sales in Receivable = 365 / Accounts Receivable Turnover = 365 / 5.4 = 67.59 Days

6.Inventory Turnover = Cost of goods sold / Avg. Inventory = 854100 / ( 321200 + 248200 ) / 2 = 854100 / 284700 = 3

7.Number of Days sales in Inventory = 365 / Inventory Turnover = 365 / 3 = 121.67 Days

8.Ratio of Fixed assets to Long Term Liabilities = 2280000 / 1900000 = 1.2

9.Ratio of Liabilities to stockholders Equity = 2489085 / 2765650 = 0.90

10.Times Interest Earned = EBIT / Interest Expense = 577500 / 152000 = 3.80

11.Asset Turnover = Sales / Avg. Total Assets = 2227230 / ( 5254735 + 4145265 ) / 2 = 2227230 / 4700000 = 0.47

12.Return on Total Assets = Net Income / Avg. Total Assets = 374400 / 4700000 = 7.97%

13.Return on Stockholders Equity = Net Income / Avg. Total Stockholders Equity

                                                  = 374400 / (2765650 + 2434350 ) / 2 = 374400 / 2600000 = 14.4%

14.Return on Common Stockholders Equity = (Net Income - preferred dividends) / Avg. Common Stockholders Equity

                                                  = ( 374400 - 11900 ) / ( 390000 + 390000 + 2035650 + 1704350 ) / 2

                                                  = 362500 / 2260000 = 16.04%

15.Earning per share = (Net Income - preferred dividends) / No. of outstanding common shares

                               = 362500 / 39000 = $9.29

16.Price Earnings Ratio = Market Price / EPS = $59 / $9.29 = 6.35

17.Dividend per share = Common stock dividend / No. of outstanding common shares = 31200 / 39000 = $0.80

18.Dividend Yield = Dividend per share / Market Price = $0.80 / 59 = 1.36%